OpenStudy (anonymous):

Sove for H =( √3-2/2 +1 ) / (√3-2/2 +2 )

OpenStudy (anonymous):

is it \[H=\frac{\frac{\sqrt{3}-2}{2}+1}{ \frac{\sqrt{3}-2}{2}+2}\]

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

lets add using this rule \[\frac{ a }{ b }+\frac{ c }{ d }=\frac{ ad +bc}{ bd }\]

OpenStudy (anonymous):

alternativley easier way is to multiply by 2 both on numerator and denominator

OpenStudy (anonymous):

can you tell me what you get when you do that

OpenStudy (anonymous):

how can I multiply it ?

OpenStudy (anonymous):

\[H=\frac{( \frac{\sqrt{3}-2}{2}+1)2 }{ (\frac{\sqrt{3}-2}{2}+2)2 }=\frac{\sqrt{3}-2+2 }{ \sqrt{3}-2+4 }=\frac{ \sqrt{3} }{ \sqrt{3}+2 }=\frac{ 3 }{ 3+2\sqrt{3} }\]

OpenStudy (anonymous):